# Consumption Partial Insurance

### v4

@@ C.6 Additional results
Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption data from the CEX. As such, this table accompanies table 3 in the main text, which reports results using consumption internally available in the PSID. In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data directly available in the PSID, regardless of whether we target only second- or also higher-order moments. This echoes our earlier discussion about the imputation inflating the pass-through of shocks. We also find that the transmission parameter of transitory shocks $\psi^{(1)}$ is negative in all cases. Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude ($-0.418$) and statistically significant. This is counterintuitive and raises further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved. We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function. However, $\phi^{(2)}<0$ in that case with a magnitude very similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant). This is, once again, indicative of the asymmetric pass-through of negative versus positive permanent shocks. Similar to our baseline, we also find $\psi^{(2)}>0$.
@@
Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution. Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted. This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^7] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets. Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, possess more insurance than the poor.
@@
Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education. As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification. The odd numbered columns present parameter estimates when only second-order moments are targeted. $\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across the groups. This is similar to what we found in the quadratic specification. However, as soon as we target up to fourth-order moments in the linear specification, shown in the even numbered columns, this monotonic pattern is lost. Transitory shocks transmit strongly in group 1, and with a negative coefficient.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary by wealth and education. The underlying income moments also vary by wealth and education (results not shown for brevity). We do not report results when up to third-order moments of income and consumption are targeted in the estimation; these are available upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies the quadratic specification estimates in table 6 in the paper.
Appendix D Welfare costs details
Table D.1 presents the parameter estimates and model fit of the quantitative model for income (mixture of normal distributions). Table D.2 calculates the welfare costs allowing for heterogeneity in income risk and consumption transmission by age, while table D.3 allows for heterogeneity by wealth. Table D.4 reports average welfare costs for different values of the utility coefficient $\gamma$.
@@ Appendix D Welfare costs details
Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$ and low (panel A) versus high wealth (panel B). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999–2019. We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text. However, the consumption parameters can be inferred from table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample. The first row presents the average costs across all households in the wealth group. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.

### v5

@@ C.6 Additional results
Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption data from the CEX. As such, this table accompanies table 3 in the main text, which reports results using consumption internally available in the PSID. In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data internally available in the PSID, regardless of whether we target only second- or also higher-order moments. This echoes our earlier discussion about the imputation inflating the pass-through of shocks. We also find that the transmission parameter of transitory shocks $\psi^{(1)}$ is negative in all cases. Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude and negative ($-0.418$). This is counterintuitive, raising further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved. We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function. $\phi^{(2)}<0$ in that case, with a magnitude similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant). Once again, this is indicative of the asymmetric pass-through of negative versus positive permanent shocks. Similar to our baseline, we also find $\psi^{(2)}>0$.
@@
Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution, respectively. Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted. This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^7] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets. Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, have more insurance against permanent income shocks than the poor.
@@
Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education. As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification. The odd-numbered columns present parameter estimates when only second-order moments are targeted. $\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across groups. This is similar to what we found in the quadratic specification. However, as soon as we target up to fourth-order moments in the linear specification, shown in the even-numbered columns, this monotonic pattern is lost. Transitory shocks transmit strongly in group 1, and with a negative coefficient.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary by wealth and education. The underlying income moments also vary by wealth and education (results not shown for brevity). We do not report results when up to third-order moments of income and consumption are targeted, but these are available upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling). All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. This table accompanies the quadratic specification estimates in table 6 in the paper.
Appendix D Some welfare costs of tail income risk
Given our empirical, quasi reduced-form approach to measuring partial insurance, we calculate the welfare costs of income risk outside of a fully specified structural model. Specifically, we draw a sequence of income shocks over the lifetime, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare based on those consumption responses. Our focus is on the costs of non-Gaussian relative to (counterfactual) Gaussian income risk, and we are particularly interested in contrasting the implications of linear and nonlinear transmission.
There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2). We assume $\mu_{1}<0$ for identification. Table D.1 reports the parameter estimates and the (excellent) model fit. In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it},$
These reflect the linear and quadratic consumption functions (4) and (5) in the paper, with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data. We simulate consumption over a working life of 35 years, starting at age 30.888We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample. In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households.
@@ Appendix D Some welfare costs of tail income risk
In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively. The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of linear transmission. Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is higher relative to a Gaussian density.
Table D.2 averages $\alpha_{i}$ across households, for $\gamma=2$. The first row presents the unconditional average cost across all households. The following rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. We also count the fraction of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime. Two main points stand out.
Table D.2: Welfare costs of non-Gaussian relative to Gaussian income risk
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
All households 100% -0.81 -0.88 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
First, the unconditional cost is small but negative in both the first (linear) and second (nonlinear) scenarios, suggesting households may in fact prefer tail over Gaussian risk. While this may seem counterintuitive, table D.1 shows that tail risk entails a small probability (about 5%) of shocks from a distribution with negative mean and large variance, and a high probability of shocks from one with a small positive mean and very low variance. Most households never actually experience shocks from the former; instead, they face the latter, whose variance is much smaller than that of Gaussian risk. Counterfactually, Gaussian risk implies larger variability for many, so households are on average willing to pay to move to the non-Gaussian setting. However, the average cost turns positive in scenario 3, which evaluates the consumption loss of tail risk and the bias from a misspecified consumption function. That is, neglecting non-Gaussian features and nonlinear transmission leads to an overestimation of consumer welfare by 0.66% of lifetime consumption.
Second, the welfare costs become substantial for households who actually experience large negative shocks at least once in life – those drawing from the first distribution in the mixture. Households facing a permanent shock at least $3\sigma$ below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once) give up 4.8% of lifetime consumption to avoid tail risk when risk is transmitted linearly (column 1, second row). This rises to 6.39% in the presence of nonlinear transmission (column 2, second row), as bad shocks transmit more strongly in this case. For worse events, costs increase further to almost 1/8 or 11.89% of lifetime consumption among households with at least one $5\sigma$ shock below the mean (column 2, fourth row; $\min\{\zeta_{it}\}<-0.95$; 17.07% of households experience this at least once), and are higher still in scenario 3 where we evaluate the welfare cost of tail risk and the bias from a misspecified consumption function. They are also (much) higher under a counterfactual zero risk scenario, like that explored in Guvenen et al. (2024).
These findings recur when we account for heterogeneity in income risk, consumption transmission, or wealth. Table D.3 calculates the welfare costs allowing for age-specific risk and consumption transmission; the costs are in the same ballpark as the baseline described above. Table D.4 further calculates the welfare costs by wealth. The least wealthy have costs almost twice as large as the baseline, while the wealthier have smaller costs. This gradient is explained by the different degrees of consumption insurance between the two groups. As the likelihood that a borrowing constraint binds depends on wealth and age, the wealth- and age-specific results are indirectly reflective also of such constraints. Without a structural model, we cannot allow households to switch wealth groups endogenously (e.g., in response to an income shock), yet the baseline costs (table D.2) lie in the middle of the wealth-specific costs (table D.4), so exogenous wealth is unlikely to have caused a serious bias in our baseline calculations. The welfare estimates are also robust to small variations in $\gamma$ (table D.5).
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$ and low (panel A) versus high wealth (panel B). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999–2019. We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text. The consumption parameters can be proxied by table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample. The first row presents the average costs across all households in the wealth group. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.